OpenStudy (anonymous):

at track practice the team runs twice around the 440 yard track does 50 jumping jacks and runs five 50-yard sprints how far does the team have to run

OpenStudy (anonymous):

440*2 + 5*50

OpenStudy (anonymous):

in addition, can you please help me with t\[\sqrt{?}\]161

OpenStudy (anonymous):

=880+250=1130 yards (too many)

OpenStudy (anonymous):

twice 440= 2*440 five 50's are 5*50 jumping jackes aren't running 880+250=930

OpenStudy (anonymous):

mean 1130 %0

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

what did you ask about \[\sqrt{?}\]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

t ( that sign ) 16=80